# Adjacency spectral radius and $H$-factors in 1-binding graphs

### v1

@@ 1 Introduction
For vertex-disjoint graphs $G_{1}$ and $G_{2}$, the graph with vertex set $V(G_{1})\cup V(G_{2})$ and edge set $E(G_{1})\cup E(G_{2})$ is called the union of $G_{1}$ and $G_{2}$, which is denoted by $G_{1}\cup G_{2}$. For any integer $k\geq 1$, let $kG$ denote the disjoint union of $k$ copies of $G$. The join of $G_{1}$ and $G_{2}$ is denoted by $G_{1}\vee G_{2}$, which is obtained from $G_{1}\cup G_{2}$ by joining each vertex of $G_{1}$ with each vertex of $G_{2}$ by an edge.
@@
A graph $G$ is said to be $r$-binding if $bind(G)\geq r$, where $r$ is a positive real number. We refer the reader to [3, 4, 5, 6, 7] for some properties on binding numbers.
Let $G$ be a graph with $V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$. The adjacency matrix of $G$ is the real symmetric matrix $A(G)=(a_{ij})_{n\times n}$, where $a_{ij}=1$ if $v_{i}v_{j}\in E(G)$ and $a_{ij}=0$ if $v_{i}v_{j}\notin E(G)$. The largest eigenvalue of $A(G)$, denoted by $\rho(G)$, is called the adjacency spectral radius of $G$. For some results on spectral radius in graphs, we refer the reader to [8, 9, 10, 11, 12, 13, 14, 15, 16].
@@ 2 Some preliminaries
Lemma 2.2. Let $\sum\limits_{i=1}^{t}n_{i}=n-s$ with $s\geq 1$. If $n_{t}\geq n_{t-1}\geq\cdots\geq n_{1}\geq 1$, $n_{t-1}\geq p\geq 1$ and $n_{t}<n-s-t-p+2$, then
@@
$\geq$ 00 0
for $1\leq i\leq t$. This leads to $x_{t}\geq x_{i}$ for $1\leq i\leq t$. Write $G^{\prime}=K_{s}\vee(K_{n-s-t-p+2}\cup K_{p}\cup(t-2)K_{1})$. Then it follows from $p\leq n_{t-1}\leq n_{t}<n-s-t-p+2$, $n_{j}\geq 1$ for $21\leq j\leq t-$ 2 and $x_{t}\geq x_{i}$ for $1\leq i\leq t$ that
@@
that is, $\rho(K_{s}\vee(K_{n_{t}}\cup K_{n_{t-1}}\cup\cdots\cup K_{n_{1}}))<\rho(K_{s}\vee(K_{n-s-t-p+2}\cup K_{p}\cup(t-2)K_{1}))$. This completes the proof of Lemma 2.2. $\Box$
@@
Proof of Theorem 1.2. Suppose that a connected 1-binding graph $G$ has no $H$-factor. Using Theorem 1.1, we conclude $\omega(G-S)\geq|S|+2$ for some nonempty subset $S$ of $V(G)$. Let $|S|=s$ and $\omega(G-S)=q$. Then $q\geq s+2$. Let $C_{1},C_{2},\ldots,C_{q-1},C_{q}$ denote the $q$ components in $G-S$ and $|C_{i}|=n_{i}$ for $1\leq i\leq q$. Without loss of generality, we let $n_{q}\geq n_{q-1}\geq\cdots\geq n_{2}\geq n_{1}\geq 1$. We are to verify $n_{s+1}\geq 2$. Otherwise, if $n_{i}=1$ for $11\leq i\leq s+$ 1, then we let $X=V(C_{1})\cup V(C_{2})\cup\cdots\cup V(C_{s+1})$. Obviously, $N_{G}(X)\subseteq S$ and $|X|=s+1$. Thus, we deduce

### v2

@@ 1 Introduction
For vertex-disjoint graphs $G_{1}$ and $G_{2}$, the graph with vertex set $V(G_{1})\cup V(G_{2})$ and edge set $E(G_{1})\cup E(G_{2})$ is called the union of $G_{1}$ and $G_{2}$, which is denoted by $G_{1}\cup G_{2}$. For any integer $k\geq 1$, let kGkG denote the disjoint union of $k$ copies of $G$. The join of $G_{1}$ and $G_{2}$ is denoted by $G_{1}\vee G_{2}$, which is obtained from $G_{1}\cup G_{2}$ by joining each vertex of $G_{1}$ with each vertex of $G_{2}$ by an edge.
@@
A graph $G$ is said to be $r$-binding if $bind(G)\geq r$, where $r$ is a positive real number. We refer the reader to [3, 4, 5, 6, 7, 8] for some properties on binding numbers.
Let $G$ be a graph with $V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$. The adjacency matrix of $G$ is the real symmetric matrix $A(G)=(a_{ij})_{n\times n}$, where $a_{ij}=1$ if $v_{i}v_{j}\in E(G)$ and $a_{ij}=0$ if $v_{i}v_{j}\notin E(G)$. The largest eigenvalue of $A(G)$, denoted by $\rho(G)$, is called the adjacency spectral radius of $G$. For some results on spectral radius in graphs, we refer the reader to [9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
@@ 2 Some preliminaries
Lemma $2.$ Let $\sum\limits_{i=1}^{t}n_{i}=n-s$ with $s\geq 1$. If $n_{t}\geq n_{t-1}\geq\cdots\geq n_{1}\geq 1$, $n_{t-1}\geq p\geq 1$ and $n_{t}<n-s-t-p+2$, then
@@
$\geq$ $0$
for $1\leq i\leq t$. This leads to $x_{t}\geq x_{i}$ for $1\leq i\leq t$. Write $G^{\prime}=K_{s}\vee(K_{n-s-t-p+2}\cup K_{p}\cup(t-2)K_{1})$. Then it follows from $p\leq n_{t-1}\leq n_{t}<n-s-t-p+2$, $n_{j}\geq 1$ for $1\leq j\leq t-2$ and $x_{t}\geq x_{i}$ for $1\leq i\leq t$ that
@@
that is, $\rho(K_{s}\vee(K_{n_{t}}\cup K_{n_{t-1}}\cup\cdots\cup K_{n_{1}}))<\rho(K_{s}\vee(K_{n-s-t-p+2}\cup K_{p}\cup(t-2)K_{1}))$. This completes the proof of Lemma $2.$ $\Box$
@@
Proof of Theorem 1.2. Suppose that a connected 1-binding graph $G$ has no $H$-factor. Using Theorem 1.1, we conclude $\omega(G-S)\geq|S|+2$ for some nonempty subset $S$ of $V(G)$. Let $|S|=s$ and $\omega(G-S)=q$. Then $q\geq s+2$. Let $C_{1},C_{2},\ldots,C_{q-1},C_{q}$ denote the $q$ components in GSG-S and $|C_{i}|=n_{i}$ for $1\leq i\leq q$. Without loss of generality, we let $n_{q}\geq n_{q-1}\geq\cdots\geq n_{2}\geq n_{1}\geq 1$. We are to verify $n_{s+1}\geq 2$. Otherwise, if $n_{i}=1$ for $1\leq i\leq s+1$, then we let $X=V(C_{1})\cup V(C_{2})\cup\cdots\cup V(C_{s+1})$. Obviously, $N_{G}(X)\subseteq S$ and $|X|=s+1$. Thus, we deduce